OpenStudy (anonymous):

Find the exact value of the expression. Do not use a calculator. If tan θ = 9, find the exact value of cot (pi/2 - θ)

OpenStudy (anonymous):

A. 8 B. 10 C. 1/9 D. 9

OpenStudy (anonymous):

don't be fooled, it is the cofunction identity \[\cot(\frac{\pi}{2}-\theta)=\tan(\theta)\]

OpenStudy (anonymous):

therefore if \(\tan(\theta)=9\) you know \(\cot(\frac{\pi}{2}-\theta)=9\) also

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

its d) 9

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

too simple